At what fraction of the speed of light does a particle travel if it's kinetic energy is twice it's rest mass


1
Expert's answer
2021-03-01T12:38:54-0500

The total energy of the particle is "E_{tot} = E_{kin}+E_{rest}=2E_{rest}+E_{rest}=3E_{rest}".

At the same time the special relativity gives us the relations "E_{tot.} = \frac{mc^2}{\sqrt{1-v^2/c^2}}", "E_{rest}=mc^2".

From this we deduce "\sqrt{1-v^2/c^2} = 1/3, v^2/c^2= 8/9" and thus "\frac{v}{c}=\frac{2\sqrt{2}}{3}".